A mobile banking app speeds up the check deposit process by two days Bank charges $.20 for each check deposited by using the app Funds can be invested at 12% annually What is the average check amount that would justify using this mobile app?

$300

$100

$400

$200
Answer #1
Correct option is: $300
Workings:
Let the average amount be X
X * 12% * 2 / 360 = 0.2
X * 0.24 = 0.2 x 360
0.24 X = 72
X = 72 / 0.24
X = 300
Therefore, the average check amount to justify the mobile app should be $300 or more